Shift Summary, 10/11/99 Day Shift

Yee Bob Hsiung (hsiung@fnal.gov)
Mon, 11 Oct 1999 15:30:06 -0500

Shift Summary for Shift, 10/11/99 Day

On Shift: S. Taegar/BobH

Hours of beam available at all intensities: 0:00
Hours of beam available at nominal intensity: 0:00
Hours of E799 data (special runs not included): 0:00
Hours of special runs (laser scans, muon runs...): 0:00

Accounting of lost time: TeVatron cool down of C1 cryo string.
TeVatron checkout.
KTeV interlocks.

========================================================================

Accesses during shift: Search&secure twice due to interlock dropped
(reason unknown - suspect spiders/spiderweb
near interlock door #, the backdoor on the
west side next to neutrino berm)

========================================================================

Overview of shift: Waiting patiently for beam, Cosmic runs.
Brainstorming the wierd L3 accept fraction problem
(CsI TE quadrant missing in 2e-nclus) in run 14794.
MCR investigate our interlock problem.

========================================================================

Description of Shift:

run spill time comment
-----------------------------------------------------------------

14815 8:30 Cleaned ktev3 drive 2d3 due to cleaning light on.
10:45 CsI TE quadrant drops out of laser illumination.
Pipeline reset/kfrends have no effect.
drt_fresh_start
14816 fixed laser illum.
11:30 UPCR01 mounted on ktev3.
11:55 stop run to install watchl3.exe to check
L3 accept fraction.
14817 12:05 cosmic run again.
12:55 MCR called to come Search&secure.
Our interlock dropped with some unknown reason.
13:40 Search&secure done.
A suspicious spiderweb found near door #2.
13:55 BobT walks-in and inform us interlock dropped
again. Call MCR to investigate. They will send
a interlock tech with a can of raid to
investigate.
14:48 MCR e-log: Interlocks reports killing
"lots of spiders" at KTev and believes KTev is
ready to be resecured. - m.m
15:13 Finished Search&secure again.
TeVatron checkout completed, MCR is trying
to establish beam in Main Injector.

========================================================================
Items to Follow up on Next Shift:

When beam comes do muon run, then pedestal run, then E799 data taking
with NM2SR - neg and NM4AN - neg pol.